the formula for the a traffic light remans yellow is t=1/8s+1 where t is time in seconds and s is the speed limit in miles per hour. Solve for s
Answer by venugopalramana(3286)

the formula for the a traffic light remans yellow is t=1/8s+1
T-1=S/8
S=8(T-1)=8T-8
where t is time in seconds and s is the speed limit in miles per hour. Solve for s